I have variables on both sides: 7p-12 = 3p
Answer by checkley75(3666)

7p-12=3p subtract -3p from both sides.
7p-3p-12=3p-3p
4p-12=0 add 12 to both sides.
4p=12 divide both sides by 4.
p=12/4 reduce the fraction.
p=3
proof
7^3-12=3*3
21-12=9
9=9